Calculate the equlibrium quantity and price if the quantity supplied can be represented by equation Qs=18000 +0. 2P and quantity demanded can be represented by the Equation Qd=2400-0.1P


1
Expert's answer
2021-04-06T18:42:21-0400

The market equilibrium occurs when "Q_d=Q_s" (it seems that there is a typo in supply function because in this case we get the negative equilibrium price):


"2400-0.1P_E=1800+0.2P_E,""0.3P_E=600,""P_E=\$2000."

Then, we can substitute "P_E" into the demand function and find the equilibrium quantity:


"Q_E=2400-0.1\cdot\$2000=2200."